OpenStudy (anonymous):

do not solve, help put into equation format please Bess and Nancy work together on one job for 8 hours. The next morning, it takes Nancy 4 hours to finish the rest of the job. If it would have taken Bess 12 hours to do the job alone, how long would it take Nancy alone?

OpenStudy (anonymous):

Convert ot hourly rates first...

OpenStudy (anonymous):

Bess takes 12 hours to do 1 job. So, in 8 hours she does (1/12)*8 = 2/3 job. Now, Nancy has done the rest of the job which is 1-2/3 = 1/3 in (8+4) = 12 hours. i.e. Nancy does 1/3 of the job in 12 hours. So, she does 1 job in 12 * 3 = 36 hours

OpenStudy (anonymous):

Sorry, I did it using the vice of common sense.